# Permutation-Invariant and Orientation-Aware Dataset Distillation for 3D Point Clouds

### v1

@@ 3.2 Permutation Invariant Distribution Matching
Random 35.52 75.17 85.31 34.57 59.96 74.06 33.96 54.84 63.13 13.85 20.42 34.75
Herding 40.07 78.02 86.86 54.38 68.45 78.84 49.49 59.79 66.85 15.67 27.66 38.67
K-center 40.07 77.59 83.22 54.38 63.03 65.26 49.49 51.39 47.82 15.67 19.81 24.00
DM 31.86 77.56 86.10 32.29 61.97 75.23 27.84 56.06 64.27 14.89 23.25 35.58
DC 43.89 74.97 86.11 52.00 66.56 75.57 49.78 60.08 64.58 18.63 24.30 35.27
Ours 44.70 84.96 87.79 55.80 72.08 80.07 50.20 63.74 68.35 17.29 31.84 43.91
Table 1: Performance comparison between three coreset selection methods [16, 2, 18], DC [29], DM [28], and the proposed method. DC, DM, and proposed method were initialized with random selection [16] for fair comparison. The best and the second best scores are highlighted in bold and underlined, respectively.
3.3 Orientation Optimization
3D models usually exhibit different orientations with each other. Therefore, while generating optimal 3D models in the synthetic dataset in terms of the geometric shapes, we also estimate their optimal orientations best representing the various orientations of 3D models in the original dataset. In practice, we employ three rotation angles of $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, around the $x$, $y$, and $z$-axes, respectively, as learnable parameters which are used to adaptively adjust the orientation of each 3D model in the synthetic dataset ${\boldsymbol{\mathcal{S}}}$ during the dataset distillation. Therefore, instead of $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (8), we minimize $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset where its 3D models are rotated toward the estimated optimal orientation.
Note that we optimize the synthetic 3D models as well as their orientations simultaneously, during the dataset distillation process. Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset. We then have $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the 3D models in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to 0. Then, for each class, a mini-batch is sampled from the original dataset $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per each class. The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with its batch size set to the number of point clouds per class (PPC). The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and the orientation parameters of its models by minimizing $\mathcal{L}_{\text{PIDM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$. This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while optimizing the orientation.
@@ 4.1 Experimental Setup
Datasets. The proposed method was evaluated on the ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively. The ScanObjectNN consists of real-world scanned objects from 15 classes, and hence a more challenging dataset due to the presence of noise, occlusions, and other artifacts commonly found in scanned 3D models.
Evaluation Metrics. Following previous methods [29, 28], we measure the classification accuracy trained on the distilled synthetic dataset. Since the training dynamics of the network can vary with each run, we ensure the fairness by training the network 10 times and report the mean accuracy across these runs. Similar to Images Per Class (IPC) in the previous image-based distillation, we evaluate the performance across different Point clouds Per Class (PPC) values. Experiments were conducted under the PPC settings of 1, 3, and 10 to analyze the impact of varying data compression levels.
Implementation Details. Each point cloud model contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15]. The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iteration by training the PointNet on the synthetic dataset and testing it on the original test dataset. Additionally, SGD was used as the optimizer for $\boldsymbol{\mathcal{S}}$, with a learning rate of 10. For optimizing the rotation angles, SGD was also used, with a learning rate of 0.5 for $\theta_{x}$ and $\theta_{z}$. Since the most datasets are aligned with the y-axis, a higher learning rate of 5 was applied for optimizing $\theta_{y}$. PointNet was optimized using the SGD optimizer with a learning rate of 0.01, and the training was conducted for 500 iterations. Also, note that we only use the distribution of the features before the max pooling to compute the PIDM loss to retain fine-grained structural information. To further emphasize the most dominant features after sorting, the loss computed from the largest magnitude values is also used.
@@ 4.2 Performance Comparison
We compare the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18]. We additionally compare DC [29] and DM [28], representative distillation methods for 2D images, to provide a broader perspective on dataset distillation. We also compare PCC [27], the only existing 3D point clouds dataset distillation method.
Table 1 compares the quantitative performance of the proposed method with that of the existing methodologies. DM and DC were initialized with random selection [16]. The synthetic datasets were optimized using PointNet [14], and classification performance was evaluated on PointNet to demonstrate their validity. Among the coreset selection methods, Herding [2, 1] achieves the best overall performance. When PPC is set to 1, Herding and K-Center [18] yield identical results since K-Center selects the first data point using the same algorithm to Herding. As PPC increases, Herding consistently outperforms other coreset selection methods, reflecting its effectiveness. However, since coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions.
DM [28] matches the distribution of the global features extracted by PointNet. However, PointNet applies the global max pooling before classification, retaining only the most dominant signal per channel while discarding substantial information. This loss of information weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset. DC [29], on the other hand, tends to fit the specific network used for distillation due to its structure designed for learning the network. However, PointNet [14], which is used for distillation, is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients. As a result, it consistently underperforms compared to the proposed method. Furthermore, as the PPC increases, the optimization process involves a larger number of point clouds, making the gradient alignment increasingly unstable and leading to reduced performance.
The proposed method outperforms both the coreset selection and existing distillation techniques. When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement. As the PPC increases, the method more effectively matches the feature distributions significantly improving the performance.
Method DM DC Ours DM DC Ours DM DC Ours DM DC Ours
PointNet++ 34.97 51.49 82.44 14.14 46.88 68.45 20.26 34.30 53.20 15.45 11.57 19.70
DGCNN 56.60 62.58 79.01 51.65 52.30 66.86 13.77 37.98 52.81 17.72 12.49 19.02
PointConv 33.41 39.96 56.57 20.00 37.90 51.32 17.35 33.99 47.41 15.16 14.20 16.82
PT 48.52 61.81 77.62 26.70 47.54 61.95 34.15 42.62 55.48 17.06 14.89 21.87
Table 2: Comparison of cross-architecture generalization performance of DC [29], DM [28], and the proposed method, evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
@@
Table 3: Performance comparison across different initialization strategies including uniform noise, random selection [16], Herding [2], and K-center [18]. We averaged the accuracies of thre cases of PPC, 1, 3, and 10. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
4.2.1 Cross-Architecture Generalization
We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31]. Unlike PointNet, the architectures used for generalization performance comparison follow a hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from PointNet. This difference makes it essential for a distilled dataset to retain rich and transferable features that generalize well across architectures.
Since the use of symmetric functions like global max pooling causes significant information loss, DM fails to preserve the original feature distribution and instead learns only the features relevant to PointNet. This limitation prevents the synthetic dataset from capturing diverse structural details, making it highly specific to PointNet. As a result, when evaluated on hierarchical models, DM performs poorly, as it lacks the necessary information for effective generalization. DC, while performing better than DM, also suffers from key limitations. Its gradient matching approach forces the synthetic dataset to align with whole gradients of PointNet parameters, which makes it more effective within the same architecture but limits its adaptability to other models. Also, the bi-level optimization process further aggravates this issue by overfitting the synthetic dataset to the specific training conditions of PointNet. Consequently, when applied to other architectures, DC struggles to generalize effectively.
The proposed method addresses these limitations by direct feature distribution matching with randomly initialized network instead of bi-level optimization. By comparing feature maps with richer information, the synthetic dataset captures features relevant to PointNet while following the original data distribution more closely. This broader feature matching strategy improves performance across different architectures. Unlike DC, which suffers from overfitting, and DM, which loses critical information, the proposed approach adapts more effectively, showing strong generalization across various backbone networks.
@@ 4.2.1 Cross-Architecture Generalization
Table 4: Comparison of training time (in hours) of DC, DM, and the proposed method. Experiments were conducted at PPC 3. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
@@
w/o sorting 75.01 59.96 53.91 20.20
w/ sorting 84.96 72.08 63.74 31.57
Table 5: Ablation study of the proposed sorting method. Experiments were performed at PPC 3. “Random” denotes the results of the random selection [16]. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
@@ 4.2.2 Initialization Strategies
Table 3 presents the classification performance on PointNet for DC [29], PCC [27] and DM [28] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18], compared to the proposed method. Note that the results of PCC are obtained from our implementation, where PCC refers to performing DC on a dataset initialized using Herding.
When initialized with noise, DC fails to effectively optimize the synthetic dataset, as the lack of geometric information disrupts the gradient computation. As a result, the original dataset fails to guide the training process along a meaningful path, leading to poor convergence. In contrast, structured initializations like random selection, Herding, or K-Center improve performance, with Herding performing best by selecting the most representative samples. Unlike DC, DM does not include network training process, making it less sensitive to noise initialization and allowing for better performance in such cases. However, when initialized with random selection, K-Center, or Herding, DM consistently underperforms DC due to its reliance on feature representations that discard significant information.
Unlike DC or DM, whose dataset distillation algorithms are unstable, the proposed method achieves comparable performance even when initialized from uniform noise. Furthermore, the proposed method significantly outperforms other approaches under the heuristic initialization strategy, such as random selection. This demonstrates that the proposed method is robust to initialization and effectively resolves the instability issues inherent in previous dataset distillation approaches.
@@
Table 6: Ablation study of optimizing the synthetic dataset and orientation. Experiments were performed at PPC 3. The baseline (first row) shows the performance without optimization which is identical to that of the random selection.
4.2.3 Training Time Comparison
Table 4 compares the training time between the proposed and previous methods. DM [28] achieves the fastest training time as it uses only a small subset of the overall features. However, this limited feature usage causes many useful features to be ignored, leading to performance degradation. In contrast, DC [29] incurs significantly higher computational cost as it aligns gradients across all parameters. Notably, for the largest dataset, ShapeNet [4], DC requires over 7 hours of training, whereas DM completes training in just 9 minutes (0.15 hours), making DC more than 50 times slower. This latency is even more pronounced in point cloud datasets, as networks designed for point cloud processing, such as PointNet [14], have higher computational complexity than those used in image dataset distillation. Since our algorithm is based on distribution matching approach, it requires only a marginal increase in training time compared to the DM algorithm, yet achieves a significant improvement in performance.
Layer 1 Layer 2 Layer 3 PointNet
_ _ 61.47
_ _ 62.22
_ _ 62.72
Table 7: Ablation study of the impact of different PointNet layers. The performance was averaged across the four models of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. Experiments were performed at PPC 3.
4.3 Visualization
Figure 3 presents a visualization of the distilled synthetic dataset generated by previous methods and the proposed method. DC [29], which relies on gradient matching, struggles to effectively match the gradients of the large PointNet [14] network. As a result, it fails to deviate significantly from the initialized point cloud model, leading to minimal movement in certain points while introducing noise. DM [28], on the other hand, performs feature distribution matching after the global max pooling operation in PointNet, which loses critical information. As a result, similar to DC, DM preserves the original shape while only shifting specific points, failing to capture meaningful structural changes. In contrast, the proposed method successfully preserves the semantic structure of each class while selectively learning essential features. Notably, in the airplane class, the wing structure appears distinct compared to other methods. Additionally, in certain categories, such as guitar, both the shape and orientation are changed, demonstrating that the proposed method functions as intended. These results highlight the effectiveness of our approach in producing a cleaner, noise-free synthetic dataset with superior visual quality compared to previous methods.
4.4 Ablation Study
4.4.1 Effect of Permutation Invariant Distribution Matching Loss
Table 5 presents an ablation study on the impact of the proposed sorting mechanism on classification performance using PointNet [14]. Without the sorting mechanism (w/o $\mathcal{F}$), the unordered nature of point clouds restricts effective feature matching, resulting in performance that is nearly identical to that of random selection [16], indicating that the model has not been properly optimized. Otherwise, when the sorting mechanism is applied (w/ $\mathcal{F}$), accuracy improves across all datasets, as the sorted feature provide a consistent and structured representation that facilitates more effective feature matching during optimization. This result confirms that addressing the inherent permutation-invariance of point clouds through sorting enables the network to align corresponding features more effectively, resulting in a substantial improvement in classification performance.
4.4.2 Effect of Orientation Optimization
@@
Table 6 presents an ablation study to evaluate the impact of optimizing the synthetic dataset and rotation angles independently and jointly, using the proposed PIDM loss. We evaluated on dataset groups categorized into aligned, mixed, and rotated to analyze the effectiveness of the proposed orientation optimization. The aligned group consists of datasets where objects maintain a consistent orientation, including ModelNet10 and ShapeNet [4]. The mixed group includes datasets where only certain classes exhibit rotation variations, represented by ModelNet40 [24]. The rotated group comprises datasets in which objects exhibit arbitrary rotations across all classes, such as ScanObjectNN [19].
In the second row, we present the results where only the rotation angle is optimized without optimizing the synthetic dataset. Introducing the proposed orientation optimization method, optimized using the PIDM loss, improves performance across all groups compared to the non-optimized setting. The improvement becomes more noticeable as the class with rotations becomes more dominant in the dataset, indicating the importance of orientation optimization in determining the optimal direction for data with arbitrary rotations. The addition of synthetic dataset optimization further boosts performance, with the best results achieved when both optimizations are applied jointly. Specifically, the accuracy on the rotated dataset increases significantly, showing that jointly optimizing the synthetic dataset and rotation angles with the PIDM loss helps the model better adapt to rotation variations.
4.4.3 Effect of Feature Extraction Layers in PointNet
Table 7 presents an ablation study examining how classification performance on PointNet [14] varies when using feature maps from different layers. PointNet consists of three layers before the pooling operation, which produce feature maps with 64, 128, and 1024 channels, respectively. The results show that deeper feature extraction layers generally lead to better performance. This performance improvement can be attributed to the increased number of channels, which allows the feature map to capture more detailed and diverse information. A higher-dimensional representation provides greater capacity to maintain essential features from the original dataset, enabling the network to better approximate the underlying distribution.

### v2

@@ 3.3 Estimation of Optimal Rotations
Random 35.5±$\pm 4.7$ 75.2±$\pm 1.7$ 85.3±$\pm 1.1$ 34.6±$\pm 1.8$ 60.0±$\pm 1.32$ 74.1±$\pm 0.4$ 34.0±$\pm 3.0$ 54.8±$\pm 1.5$ 63.1±$\pm 1.0$ 13.9±$\pm 1.4$ 20.4±$\pm 1.3$ 34.8±$\pm 1.1$
Herding 40.1 ±$\pm 5.2$ 78.0 ±$\pm 1.3$ 86.9±$\pm 0.6$ 54.4 ±$\pm 2.0$ 68.5 ±$\pm 1.0$ 78.8 ±$\pm 0.4$ 49.5 ±$\pm 2.3$ 59.8 ±$\pm 0.9$ 66.9 ±$\pm 0.5$ 15.7 ±$\pm 1.7$ 27.7 ±$\pm 1.4$ 38.7 ±$\pm 1.7$
K-center 40.1 ±$\pm 5.2$ 77.6±$\pm 1.9$ 83.2±$\pm 1.4$ 54.4±$\pm 2.0$ 63.0 ±$\pm 2.7$ 65.3 ±$\pm 1.1$ 49.5 ±$\pm 2.3$ 51.4 ±$\pm 1.7$ 47.8 ±$\pm 0.7$ 15.7 ±$\pm 1.7$ 19.8 ±$\pm 0.8$ 24.0 ±$\pm 1.0$
DM 31.9±$\pm 4.3$ 77.6±$\pm 1.6$ 86.1±$\pm 0.8$ 32.3±$\pm 4.8$ 62.0±$\pm 1.9$ 75.2±$\pm 0.5$ 27.8±$\pm 4.0$ 56.1±$\pm 1.2$ 64.3±$\pm 0.7$ 14.9±$\pm 2.0$ 23.3±$\pm 1.5$ 35.6±$\pm 1.4$
DC 43.9±$\pm 5.5$ 75.0±$\pm 2.4$ 86.1±$\pm 1.1$ 52.0±$\pm 1.7$ 66.6±$\pm 1.3$ 75.6±$\pm 0.5$ 49.8±$\pm 1.3$ 60.1±$\pm 1.4$ 64.6±$\pm 0.7$ 18.6±$\pm\textbf{1.6}$ 24.3±$\pm 2.5$ 35.3±$\pm 1.2$
Ours 44.7 ±$\pm\textbf{6.1}$ 84.4 ±$\pm\textbf{1.2}$ 87.8 ±$\pm\textbf{1.0}$ 55.8 ±$\pm\textbf{1.5}$ 72.1 ±$\pm\textbf{0.8}$ 80.1 ±$\pm\textbf{0.4}$ 50.2 ±$\pm\textbf{1.4}$ 63.7 ±$\pm\textbf{0.7}$ 68.4 ±$\pm\textbf{0.5}$ 17.4 ±$\pm 1.5$ 31.6 ±$\pm\textbf{1.0}$ 43.9 ±$\pm\textbf{1.9}$
Table 2: Comparison of cross-architecture generalization performance evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
Method DM DC Ours DM DC Ours DM DC Ours DM DC Ours
PointNet++ 35.0±$\pm 8.3$ 51.5±$\pm 7.8$ 82.9±$\pm\textbf{1.4}$ 14.1±$\pm 2.7$ 46.9±$\pm 6.0$ 68.5±$\pm\textbf{0.5}$ 20.3±$\pm 3.7$ $\pm 3.4$ 53.2±$\pm\textbf{1.5}$ 15.5±$\pm 3.6$ 11.6±$\pm 2.2$ 22.0±$\pm\textbf{3.9}$
DGCNN 56.6±$\pm 5.3$ 62.6±$\pm 2.2$ 79.0±$\pm\textbf{2.2}$ 34.8±$\pm 3.5$ 52.3±$\pm 2.7$ 66.9±$\pm\textbf{1.1}$ 13.8±$\pm 2.2$ 38.0±$\pm 2.7$ 52.8±$\pm\textbf{1.6}$ 17.7±$\pm 2.4$ 12.5±$\pm 1.3$ 19.0±$\pm\textbf{1.5}$
PointConv 33.4±$\pm 6.2$ 40.0±$\pm 10.5$ 56.6±$\pm\textbf{4.9}$ 20.0±$\pm 5.7$ 37.9±$\pm 3.9$ 51.3±$\pm\textbf{6.2}$ 17.4±$\pm 3.0$ 34.0±$\pm 4.8$ 47.4±$\pm\textbf{2.3}$ 15.2±$\pm 1.7$ 14.2±$\pm 1.8$ 16.8±$\pm\textbf{1.5}$
PT 48.5±$\pm 6.6$ $\pm 1.6$ 77.6±$\pm\textbf{1.6}$ 26.7±$\pm 5.0$ 47.5±$\pm 2.7$ 61.6±$\pm\textbf{1.1}$ 34.2±$\pm 5.3$ 42.6±$\pm 2.5$ 55.5±$\pm\textbf{0.5}$ 17.1±$\pm 0.8$ 14.9±$\pm 0.5$ 21.9±$\pm\textbf{2.7}$
@@ 4.1 Experimental Setup
The proposed method was evaluated on the ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] datasets. The ModelNet10, ModelNet40, and ShapeNet are synthetic datasets generated from CAD models, containing 10, 40, and 55 classes, respectively. The ScanObjectNN consists of real-world scanned objects from 15 classes. Following previous methods [29, 28], we measure the classification accuracy trained on the distilled synthetic dataset. We ensure the fairness by training the network 10 times and report the mean and standard deviation of accuracy. We evaluate the performance across different PPC values of 1, 3, and 10, respectively. Each point cloud object contains 1,024 points, which is a common standard used in the 3D point clouds classification tasks [14, 15]. The dataset distillation process for the synthetic dataset was optimized for 1,500 iterations, where the performance was evaluated at every 250 iteration by training the PointNet on the synthetic dataset and testing it on the original test dataset.
@@ 4.2 Performance Comparison
4.2.1 Dataset Distillation
We compare the performance of the proposed method with that of 1) the three most representative coreset selection methods: random selection [16], Herding [2, 1], and K-Center [18], 2) two existing image dataset distillation methods of DC [29] and DM [28], and 3) the only existing 3D point clouds dataset distillation method of PCC [27]. Table 1 compares the quantitative performance of the proposed method with that of the existing methods. DM and DC were initialized with random selection. The synthetic datasets were optimized using PointNet [14], and classification performance was evaluated on PointNet to demonstrate their validity. Among the coreset selection methods, Herding achieves the best overall performance. When PPC is set to 1, Herding and K-Center yield identical results since K-Center selects the first data point using the same algorithm to Herding. As PPC increases, Herding consistently outperforms other coreset selection methods, showing its effectiveness. However, since the coreset selection methods rely only on sample selection without further optimization, they struggle to capture complex feature distributions. The features used in DM are extracted from the layer before the classifier, and they discard substantial information through the global max pooling process of PointNet. This information loss weakens the effectiveness of the distribution matching, preventing DM from accurately capturing the feature distribution of the original dataset. DC, on the other hand, aims to match the gradients of the network. However, PointNet is significantly larger than typical networks in image-based distillation tasks, making it difficult for DC to precisely align the gradients. As a result, it consistently underperforms compared to the proposed method. The proposed method outperforms both the coreset selection and existing distillation techniques. When the PPC is small, accurately capturing the original distribution is challenging, leading less noticeable improvement. As the PPC increases, the proposed method more effectively matches the feature distributions significantly improving the performance.
Table 3: Performance comparison across different initialization strategies. The accuracies of three PPC values of 1, 3, and 10 are averaged. MN10: ModelNet10. MN40: ModelNet40. SN: ShapeNet. SONN: ScanObjectNN.
@@ 4.2.1 Dataset Distillation
† DC [29] with Herding initialization corresponds to PCC [27]
Table 4: Comparison of training time (in hours) at the PPC value of 3.
@@
Table 5: Ablation study of the proposed SADM loss. Experiments were performed at the PPC value of 3. “Random” denotes the results of the random selection [16], and SA is the proposed semantic alignment.
@@
w/o SA 75.01 59.96 53.91 20.20
w/ SA 84.42 72.08 63.74 31.84
4.2.2 Cross-Architecture Generalization
We also present the cross-architecture generalization performance in Table 2, where the synthetic dataset is optimized using PointNet [14] and evaluated on different backbone networks including PointNet++[15], DGCNN[22], PointConv [23], and PT [31], respectively. Unlike PointNet, the architectures used for generalization performance comparison follow hierarchical structure similar to CNN-based networks in image processing, exhibiting significantly different characteristics from that of PointNet. This difference makes it essential for the distilled dataset to retain rich and transferable features that generalize well toward different architectures. Due to significant information loss caused by the pooling operation, DM primarily captures the features specific to the PointNet, failing to preserve the original feature distribution. Consequently, the synthetic dataset lacks diverse structural details, resulting in poor generalization performance, particularly on hierarchical models. DC, while performing better than DM, also suffers from key limitations. It matches the gradients obtained from a trained PointNet, the synthetic dataset becomes overfitted to that specific architecture and fails to generalize toward other networks. Consequently, the generalization ability of DC degrades when applied to other architectures. In contrary, the proposed method addresses these limitations by directly matching the feature distributions extracted from randomly initialized network. By comparing feature maps containing richer and less architecture-biased information, the synthetic dataset captures the features relevant to the PointNet while following the original data distribution more closely. Wherease DC suffers from the overfitting and DM loses critical information, the proposed method yields higher adaptability showing strong generalization performance across various backbone networks.
@@ 4.2.3 Initialization Strategies
Table 5 presents the classification performance on PointNet for DC [29], DM [28], and PCC [27] under different initialization strategies, including uniform noise, random selection [16], Herding [2, 1], and K-Center [18]. Note that the results of PCC are obtained by our implementation, where the PCC is equivalent to performing DC with Herding initialization. When initialized with noise, DC fails to effectively optimize the synthetic dataset because the network has far more parameters than typical networks used for image dataset distillation, making it difficult to align the gradients. Therefore, the original dataset cannot properly guide the training process, leading to poor convergence. In contrast, the structured initializations such as random selection, Herding, or K-Center improve the performance, where the Herding performs the best by selecting the most representative samples. Unlike DC, DM directly matches the feature representations rather than relying on the gradients and hence less sensitive to the initialization strategies. However, due to the information loss, DM consistently underperforms DC even with the structured initializations. In contrast, the proposed method achieves favorable performance, even when initialized from uniform noise. Furthermore, the proposed method significantly outperforms other approaches under the heuristic initialization strategy, such as random selection. This demonstrates that the proposed method is robust with diverse initialization schemes and effectively resolves the instability issues inherent in the existing dataset distillation methods.
4.2.4 Training Time
Table 5 compares the training times between the proposed method and the existing methods. DM [28] achieves the fastest training time as it uses only a small subset of the overall features, however it often ignores many useful features leading to performance degradation. In contrast, DC [29] incurs a significantly higher computational cost as it aligns the gradients across all the parameters. Notably, for the largest dataset of ShapeNet [4], DC requires over 7 hours for training 50 times slower that DM that completes the training in just 9 minutes (0.15 hours). Since the proposed method is based on the distribution matching approach, its training time is marginally increased compared to DM, yet achieves significant improvement in dataset distillation performance.
4.2.5 Qualitative Comparison of Synthetic Datasets
Figure 3 compares the resulting synthetic datasets distilled by using the proposed method and the existing methods, respectively. The first row shows the original datasets and the second row shows initialized point cloud objects. DC [29] fails to deviate significantly from the initialized objects and causes noise. Similarly, DM [28] also maintains the original shape of the initialized objects while shifting certain points only, that hinders to capture meaningful structural changes. In contrast, as shown in the bottom row, the proposed method successfully preserves the overall semantic structures of 3D objects in each class, while selectively learning essential features. For example, in the airplane class, we observe significant changes in the edges and corners, that are critical for class discrimination, especially around the wings. Also, in the guitar class, the proposed method jointly optimizes the shape and orientation of the synthesized 3D objects.
4.3 Ablation Study
4.3.1 Effect of SADM Loss
Table 5 presents the ablation study on the impact of the proposed SADM loss on the classification performance using PointNet [14]. Without preserving the semantic alignment, the unordered nature of point clouds restricts effective feature matching, resulting in nearly identical performance to that of the random selection [16]. That means the model has not been properly optimized. Otherwise, when the semantic alignment is applied, accuracy improves across all datasets, as the aligned featured provide consistent and structured representation that facilitates more effective feature matching during the optimization. This result confirms that preserving the inherent semantic alignment via feature sorting facilitates the network to exploit well-aligned corresponding features and significantly improves the classification performance.
4.3.2 Effect of Optimal Rotation Estimation
Table 6: Ablation study of the proposed optimal rotation estimation. Experiments were performed at the PPC value of 3.
@@ 4.3.2 Effect of Optimal Rotation Estimation
Table 6 presents the ablation study to evaluate the impact of joint optimization of synthetic dataset and rotation angles. We categorized the test datsets into the aligned, mixed, and rotated groups to analyze the effectiveness of the proposed rotation estimation. The aligned group consists of the datasets where the objects maintain consistent orientations, including ModelNet10 and ShapeNet [4]. The mixed group includes ModelNet40 [24] where only certain classes exhibit rotation variations. The rotated group is ScanObjectNN [19] where the objects exhibit arbitrary rotations across all classes. The baseline (first row) shows the performance without optimization which is identical to that of the random selection. In the second row, we present the results where only the orientation of the synthetic data is optimized using the proposed SADM loss, without updating the synthetic dataset. This leads to performance improvements across all the groups compared to random selection. The improvement becomes more significant as the classes with rotation variation become more dominant. Optimizing the synthetic dataset further boosts the performance, with the best results achieved when both the synthetic data and rotation angles are optimized simultaneously. In particular, the accuracy on the rotated dataset improves significantly, demonstrating that the proposed joint optimization enables the model to be resilient to rotation variations.
4.4 Limitation
While the proposed SADM loss enhances the semantic consistency by sorting the features, it does not completely align semantically meaningful regions across different point cloud objects. Estimating optimal rotations also improves the complexity, particularly when the datasets are already well-aligned to the canonical axes.